# Shape-constrained density estimation with Wasserstein projection

### 09 Mar 2026 v1

@@ 5. Implementation and examples
In this section, we implement monotone and log-concave density estimation with Wasserstein projection upon suitable discretizations, and provide examples. [^7] We refer the reader to the Github repository https://github.com/tkl-wong/wasserstein-projection-estimation for the codes and details of our implementation. Here, our purpose is to illustrate the empirical behaviors of the Wasserstein projection estimator; further work is needed to develop more efficient and scalable algorithms and, more importantly, an understanding of when estimators based on optimal transport may be more suitable for real data. [^8] In a parametric framework, it was shown recently in [35] that the Wasserstein projection is asymptotically optimal, in a suitable sense, under infinitesimal additive perturbation of the original data set.
We focus on the case $p=2$, but note that the algorithms can be modified for general $p>1$ (even $p=1$ but then uniqueness of the solution can no longer be guaranteed.). We will be working with quantile functions to utilize convexity of the Wasserstein projection. Consider a data-set represented as an empirical distribution $\mu_{n}=\sum_{i=1}^{n}\frac{1}{n}\delta_{x_{i}}$, or more generally $\mu_{n}=\sum_{i=1}^{n}w_{i}\delta_{x_{i}}$ where $w_{i}$ is the weight of $x_{i}$. Let a partition
@@
This is the case if $\Pi$ is a refinement of the uniform grid $(\frac{i}{n})_{i=0}^{n}$ (or its weighted analogue). Otherwise, we may discretize and let $y_{i}$ be the $100u_{i}\%$-quantile of $\mu_{n}$. The reason why discretization is necessary is that the sets of break points in Theorems 3.6 and 4.7 generally do not correspond to subsets of the data points and their quantile probabilities, and currently we do not have explicit characterizations of their locations.
@@ 5.1. Monotone density estimation
We will compare the Wasserstein projection estimator with Grenander’s estimator [15]. It is well known that the density of Grenander’s estimator is given by the slopes of the least concave majorant of the distribution function of $\mu_{n}$. Equivalently, the quantile function of Grenander’s estimator is the greatest convex minorant of $Q_{0}$. [^9] More precisely, we enforce the constraint that the quantile function is 0 at $u=0$. This guarantees that the estimated density is non-increasing on $\mathbb{R}_{+}$. We implement the greatest convex minorant, and hence Grenander’s estimator, using the R package isotone [12].
@@
Consider $\mu_{n}=\sum_{i=1}^{1000}\frac{1}{1000}\delta_{x_{i}}$, where $x_{i}$ is the $100v_{i}$% quantile of $\mu^{*}=0.6\mathrm{Gamma}(5,1.5)+0.4\mathrm{Gamma}(22,1.5)$ (the second slot represents the rate parameter), and $v_{i}=0.001+\frac{0.999-0.001}{1000}(i-1)$ interpolates uniformly between $0.001$ and $0.999$. These parameters are chosen for visual purposes. The quantile function of $\mu_{n}$ has two concave regions corresponding to the two peaks of the density. The Wasserstein projection $\hat{\mu}_{n}$ is approximately the $\mathcal{W}_{2}$-projection of $\mu^{*}$ onto $\mathcal{F}_{m,2}$ (see Theorem 2.14). On the other hand, Grenander’s estimator is approximately the projection with respect to the Kullback–Leibler divergence. In Figure 2, we show their quantile and density functions. Both estimators track closely the right tail of the true density. By construction, the Wasserstein projection provides a better overall fit of the empirical quantile function with respect to the $L^{2}$ distance, while Grenander’s estimator returns its greatest convex minorant. Observe that the smaller density of $\hat{\mu}_{n}$ on the left end leads to a closer fit of the quantile function around the interval $[0.5,0.8]$ (this effect is not apparent from the density plot). On the other hand, Grenander’s estimator traces $Q_{0}$ more closely around $u\approx 0.45$ and $u\approx 0.9$ where $Q_{0}$ turns upward. This shows that the two estimators lead to different tradeoffs.
@@ 5.2. Log-concave density estimation
where $\delta_{i}:=h_{i}/h_{i-1}-1>-1$ and $\Phi:(-1,\infty)\rightarrow(0,1)$ is the convex function defined by
@@
simply by the general purpose optimization R package nloptr [17]. We implement the univariate log-concave MLE using the R package logcondens [29]. It is well known that the density of the MLE is supported on the convex hull of the data. Moreover, it is piecewise log-affine and each knot must occur at one of the data points.
@@ 6. Discussion
    In Theorems 3.6 and Theorem 4.7, we showed that the estimated density is piecewise affine/log-concave. Our proof techniques do not provide further information about the number of possible break points and their locations. Improved understanding of these break points will likely suggest more precise and efficient algorithms.
@@
    Further statistical properties of the Wasserstein projection estimator. Here, finite sample properties are only investigated in Wasserstein distance when $p=2$. A natural direction is to investigate the properties of the estimated density.

### 10 Apr 2026 v2

@@ 5. Implementation and examples
In this section, we implement monotone and log-concave density estimation with Wasserstein projection upon convexity-preserving discretizations, and provide examples. [^9] We refer the reader to the Github repository https://github.com/tkl-wong/wasserstein-projection-estimation for the codes and details of our implementation. Our purpose here is simply to illustrate the empirical behaviors of the Wasserstein projection estimator; further work is needed to develop more efficient and scalable algorithms and, more importantly, an understanding of when estimators based on optimal transport may be more suitable for real data. [^10] In a parametric framework (and under suitable regularity conditions), it was shown recently in [36] that the Wasserstein projection is asymptotically optimal, in a suitable sense, under infinitesimal additive perturbation of the original data set.
We focus on the case $p=2$, but note that the algorithms can be modified for general $p\geq 1$ (when $p=1$ uniqueness of the solution can no longer be guaranteed). We will be working with quantile functions to utilize convexity of the Wasserstein projection. Consider a data-set represented as an empirical distribution $\mu_{n}=\sum_{i=1}^{n}(1/n)\delta_{x_{i}}$, or more generally $\mu_{n}=\sum_{i=1}^{n}w_{i}\delta_{x_{i}}$ where $w_{i}$ is the weight of $x_{i}$.
Let a partition
@@
This is the case if $\Pi$ is a refinement of the uniform grid $(i/n)_{i=0}^{n}$ (or its weighted analogue). Otherwise, we may discretize and let $y_{i}$ be the $100u_{i}\%$-quantile of $\mu_{n}$. The reason why discretization is necessary is that the sets of break points in Theorems 3.6 and 4.7 generally do not correspond to subsets of the data points and their quantile probabilities. Further study of these break points is a natural direction for future research.
@@ 5.1. Monotone density estimation
We will compare the Wasserstein projection estimator with Grenander’s estimator [14]. It is well known that the density of Grenander’s estimator is given by the slopes of the least concave majorant of the distribution function of $\mu_{n}$. Equivalently, the quantile function of Grenander’s estimator is the greatest convex minorant of $Q_{0}$. [^11] More precisely, we enforce the constraint that the quantile function is 0 at $u=0$. This guarantees that the estimated density is non-increasing on $\mathbb{R}_{+}$. We implement the greatest convex minorant, and hence Grenander’s estimator, using the R package fdrtool [18].
@@
Consider $\mu_{n}=\sum_{i=1}^{1000}(1/1000)\delta_{x_{i}}$, where $x_{i}$ is the $100v_{i}$% quantile of $\mu^{*}=0.6\mathrm{Gamma}(5,1.5)+0.4\mathrm{Gamma}(22,1.5)$ (the second slot represents the rate parameter) and the sequence $(v_{i})_{i=1}^{1000}$ is a uniform grid between $0.001$ and $0.999$. These parameters are chosen for visual purposes. The quantile function of $\mu_{n}$ has two concave regions corresponding to the two peaks of the density. The Wasserstein projection $\hat{\mu}_{n}$ is approximately the $\mathcal{W}_{2}$-projection of $\mu^{*}$ onto $\mathcal{F}_{m,2}$ (see Theorem 2.14). On the other hand, Grenander’s estimator is approximately the projection with respect to the Kullback–Leibler divergence. In Figure 2, we show their quantile and density functions. Both estimators track closely the right tail of the true density. By construction, the Wasserstein projection provides a better overall fit of the empirical quantile function with respect to the $L^{2}$ distance, while Grenander’s estimator returns its greatest convex minorant. Observe that the smaller density of $\hat{\mu}_{n}$ on the left end leads to a closer fit of the quantile function around the interval $[0.5,0.8]$ (this effect is not apparent from the density plot). On the other hand, Grenander’s estimator traces $Q_{0}$ more closely around $u\approx 0.45$ and $u\approx 0.9$ where $Q_{0}$ turns upward. This shows that the two estimators lead to different tradeoffs.
@@ 5.2. Log-concave density estimation
where $\delta_{i}:=h_{i}/h_{i-1}-1>-1$ and $\Phi:(-1,\infty)\rightarrow(0,\infty)$ is the convex decreasing function defined by
@@
simply by the general purpose optimization R package nloptr [17]. We implement the univariate log-concave MLE using the R package logcondens [30]. It is well known that the density of the MLE is supported on the convex hull of the data. Moreover, it is piecewise log-affine and each knot occurs at one of the data points.
@@ 6. Discussion
    In Theorems 3.6 and Theorem 4.7, we showed that the estimated density is piecewise affine/log-concave. Our proof techniques do not provide further information about the number of break points and their locations. Improved understanding of these break points will likely suggest more precise and efficient algorithms.
@@
    Further statistical properties of the Wasserstein projection estimator. Here, finite sample properties are only investigated in Wasserstein distance when $p=2$. A natural direction is to investigate convergence properties of the estimated density.